OpenStudy (anonymous):

A simple random sample of 50 adults was surveyed, and it was found that the mean amount of time that they spend surfing the Internet per day is 54.2 minutes, with a standard deviation of 14.0 minutes. What is the 99% confidence interval for the number of minutes that an adult spends surfing the Internet per da

OpenStudy (anonymous):

The confidence interval will have the form\[\left(\bar{x}-Z_{\alpha/2}\frac{\sigma}{\sqrt n},~\bar{x}+Z_{\alpha/2}\frac{\sigma}{\sqrt n}\right)\] where \(\bar{x}\) is the sample/estimated mean, \(Z_{\alpha/2}\) is the critical value for a \((1-\alpha)\times100\%\) confidence level, \(\sigma\) is the sample or population standard deviation (depending on which is known), and \(n\) is the sample size. If you're not familiar with the \(Z_{\alpha/2}\) notation, it's basically the \(z\) values that give the drawn areas below.